# SPECIAL RELATIVITY

### v9

@@ 16.5 Final Remarks
Finally, we make some remarks on another way in which our complicated problem can be solved. We know the result for Lorentz coordinatization. Now let us understand why the emission of coherent undulator radiation from the modulated electron beam in the kicked direction exist in the framework of absolute time synchronization. The physics is actually very simple, and if we treat the production of coherent undulator radiation by kicked electron beam as the aberration of light we shall see that we can understand almost everything that happens in the undulator. Our Galilean transformed electrodynamics says that by making a measurement on the coherent radiation, one can observe only radiation in the kicked direction. This is exactly analogous to the aberration (deviation of the energy transport) for light radiated from the single moving emitter in an inertial frame, Fig. 1.
@@
There is an intuitively plausible way to understanding the aberration of light from a kicked electron beam. The aberration of light in an inertial frame can be easily explained on the basis of corpuscular theory, Fig. 48. This phenomenon is fully understandable in terms of transformation of velocities between different reference frames. In the case of the transversely moving electron beam in an undulator, intuition would seem to tell us that aberration increment would be the same. This is plausible if one keeps in mind that a (undulator) radiation pulse represents a certain amount of electromagnetic energy. Energy, like mass, is a quantity that is conserved, so that a coherent X-ray pulse resembles, in many aspects, a material particles. Therefore, we should expect that group velocities of undulator radiation pulses obey the same addition theorem for particle velocities in the inertial lab frame. A closer treatment based on wave theory of light confirm this expectation.
@@
Now let us demonstrate that the Galilean transformed electrodynamics will give a deviation of the energy transport direction for radiated X-ray pulse. The modulated electron beam with finite aperture in an undulator is a kind of active medium which breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the density of the elementary source varies in the transverse direction according to the law $\rho_{e}=g(K_{\perp})\cos{({K}_{\perp}y)}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated electron beam density. From the Galilean transformation, after partial differentiation, one obtains wave equation Eq.(12). The new terms that have to be put into the field equations due to use of Galilean transformation lead to the prediction of the Doppler effect. As one of the consequences of the Doppler effect in the absolute time coordinatization, we find an angular frequency dispersion of the light waves radiated from the kicked (modulated) electron beam with finite aperture. The Doppler shift, $\Delta\omega$, of radiated light wave (in the first order approximation) is given by $\Delta\omega=K_{\perp}v_{y}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation states that radiated coherent X-ray beam with finite transverse size moves along the $y$ direction with group velocity $d\omega/dk_{y}=v_{y}$. Thus, according to correct coupling conventional particle tracking and electrodynamics, we identify the direction of X-ray beam propagation with direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
16.6 Bibliography and Notes
1. An angular kick is often an essential part of many XFEL related diagnostic or experimental procedures. The standard gain length measurement procedure in XFELs employs such kicks. Other applications include ”beam-splitting” schemes where different polarization components are separated by means of an angular kick to the modulated electron beam [61, 62].
2. In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis. All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis. However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”). The discrepancy between directions of the electron motion and wavefront normal after the kick have been discussed in the literature. One particular consequence that received attention following the [63] is the effect of the trajectory error (single kick error) on the XFEL amplification process. It was pointed out that coherent radiation is emitted towards the wavefront normal of the beam modulation. Thus, according to conventional coupling of fields and particles (which we claimed incorrect), the discrepancy between the two directions decreases the radiation efficiency [63]. Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements. However, semi-analytical studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles. The pleasant surprise is that the tolerances predicted are more stringent than they need be according to the corrected XFEL theory. This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
@@ 16.6 Bibliography and Notes
3. The fact that our theory predicts reality in a satisfactory way is well-illustrated by comparing the prediction we just made with the results of an experiment involving ”X-ray beam splitting” of a circularly-polarized XFEL pulse from the linearly-polarized XFEL background pulse, a technique used in order to maximize the degree of circular polarization. The XFEL experiments apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of motion of the kicked beam, Fig. 73. This is the only way to justify coherent radiation emission from the short undulator placed after the kicker and along the kicked direction, see Fig. 14 in [64]. These results came unexpectedly, but from a practical standpoint, the ”apparent wavefront readjusting” immediately led to the realization that the unwanted, linearly-polarized radiation background could be fully eliminated without extra-hardware.
@@
4. In existing literature a theoretical analysis of XFELs driven by an electron beam with wavefront tilt was presented in [65, 66], based on the use the usual Maxwell’s equations. In fact, the Maxwell solver was used as a part of the standard simulation code. We state that this approach is fundamentally incorrect. In ultrarelativistic asymptote a modulation wavefront tilt is absurd with the viewpoint of Maxwell’s electrodynamics. In the case of an XFEL we deal with an ultrarelativistic electron beam and within the Lorentz lab frame (i.e. within the validity of the Maxwell’s equations) the tilted modulation wavefront is at odds with the principle of relativity.

### v10

@@ 20.5 Discussion
Finally, we make some remarks on another way in which our complicated problem can be solved. We know the result for Lorentz coordinatization. Now let us understand why the emission of coherent undulator radiation from the modulated electron beam in the kicked direction exists in the framework of absolute time synchronization. The physics is actually very simple, and if we treat the production of coherent undulator radiation by kicked electron beam as the aberration of light we shall see that we can understand almost everything that happens in the undulator. Our Galilean-transformed electrodynamics says that by making a measurement of the coherent radiation, one can observe only radiation in the kicked direction. This is exactly analogous to the aberration (deviation of the energy transport) for light radiated from the single moving emitter in an inertial frame, Fig. 1.
@@
There is an intuitively plausible way to understand the aberration of light from a kicked electron beam. The aberration of light in an inertial frame can be easily explained on the basis of corpuscular theory, Fig. 39. This phenomenon is fully understandable in terms of the transformation of velocities between different reference frames. In the case of the transversely moving electron beam in an undulator, intuition would seem to tell us that the aberration increment would be the same. This is plausible if one keeps in mind that a (undulator) radiation pulse represents a certain amount of electromagnetic energy. Energy, like mass, is a quantity that is conserved, so that a coherent X-ray pulse resembles, in many aspects, material particles. Therefore, we should expect that group velocities of undulator radiation pulses obey the same addition theorem for particle velocities in the inertial lab frame. A closer treatment based on the wave theory of light confirms this expectation.
@@
Now let us demonstrate that the Galilean transformed electrodynamics will give a deviation of the energy transport direction for radiated X-ray pulse. The modulated electron beam with a finite aperture in an undulator is a kind of active medium that breaks up the radiated beam into a number of diffracted beams of plane waves. Each of these beams corresponds to one of the Fourier components into which an active medium can be resolved. Let us assume that the density of the elementary source varies in the transverse direction according to the law $\rho_{e}=g(K_{\perp})\cos{({K}_{\perp}y)}$, where $K_{\perp}$ is the wavenumber of sinusoidally space-modulated electron beam density. From the Galilean transformation, after partial differentiation, one obtains wave equation Eq.(12). The new terms that have to be put into the field equations due to the use of Galilean transformation lead to the prediction of the Doppler effect. As one of the consequences of the Doppler effect in the absolute time coordinatization, we find an angular frequency dispersion of the light waves radiated from the kicked (modulated) electron beam with a finite aperture. The Doppler shift, $\Delta\omega$, of a radiated light wave (in the first order approximation) is given by $\Delta\omega=K_{\perp}v_{y}$, where $K_{\perp}$ is the transverse component of the radiated wavenumber vector. The last equation states that a radiated coherent X-ray beam with finite transverse size moves along the $y$ direction with group velocity $d\omega/dk_{y}=v_{y}$. Thus, according to correct coupling conventional particle tracking and electrodynamics, we identify the direction of X-ray beam propagation with the direction of energy propagation, supposing the latter to transform differently from the wave normal under Galilean transformation.
@@ 20.6 Experimental Test of Coupling between the Electrodynamics and Dynamics
The fact that our theory predicts reality in a satisfactory way is well-illustrated by comparing the prediction we just made with the results of an experiment involving ”X-ray beam splitting” of a circularly-polarized XFEL pulse from the linearly-polarized XFEL background pulse, a technique used in order to maximize the degree of circular polarization. The XFEL experiments demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of motion of the kicked beam, Fig. 62. This is the only way to justify coherent radiation emission from the short undulator placed after the kicker and along the kicked direction, see Fig. 14 in [76]. These results came unexpectedly, but from a practical standpoint, the ”apparent wavefront readjusting” immediately led to the realization that the unwanted, linearly polarized radiation background could be fully eliminated without extra hardware.
@@
In existing literature a theoretical analysis of XFELs driven by an electron beam with wavefront tilt was presented in [77, 78], based on the use the Maxwell’s equations. In fact, the Maxwell solver was used as a part of the standard simulation code. We state that this approach is fundamentally incorrect. In ultrarelativistic asymptote, a modulation wavefront tilt is absurd from the viewpoint of Maxwell’s electrodynamics. In the case of an XFEL we deal with an ultrarelativistic electron beam and within the Lorentz lab frame (i.e. within the validity of Maxwell’s equations) the tilted modulation wavefront is at odds with the special relativity.
It should be noted that the results of the beam splitting experiment at LCLS confirm our correction for spontaneous undulator emission [74]. It apparently demonstrated that after a modulated electron beam is kicked on a large angle compared to the divergence of the XFEL radiation, the modulation wavefront is readjusted along the new direction of the motion of the kicked beam. Therefore, coherent radiation from the undulator placed after the kicker is emitted along the kicked direction practically without suppression (see Chapter 19 for more detail). In the framework of the conventional theory, there is a second outstanding puzzle concerning the beam-splitting experiment at the LCLS. In accordance with conventional undulator radiation theory, if the modulated electron beam is at perfect resonance without the kick, then after the kick the same modulated beam must be at perfect resonance in the velocity direction. However, experimental results clearly show that when the kick is introduced there is a redshift in the resonance wavelength. The maximum power of the coherent radiation is reached when the undulator is detuned to be resonant to the lower longitudinal velocity after the kick [74]. It should be remarked that any linear superposition of a given radiation field from single electrons conserves single-particle characteristics like parametric dependence on undulator parameters and polarization. Consider a modulated electron beam kicked by a weak dipole field before entering a downstream undulator. Radiation fields generated by this beam can be seen as a linear superposition of fields from individual electrons. Now experimental results clearly show that there is a redshift in the resonance wavelength for coherent undulator radiation when the kick is introduced. It follows that the undulator radiation from the single electron has a redshift when the kick is introduced as well. This argument suggests that the results of the beam splitting experiment in reference [74] confirm our correction for spontaneous undulator emission.
20.7 Myth about the Reality of the Modulation Wavefront Tilt
In existing XFEL theory the wavefront tilt is considered as real. However, there is a common mistake made in accelerator physics connected with the wavefront tilt. In the ultrarelativistic domain the wavefront tilt has no exact objective meaning. The angle of wavefront tilt depends on the choice of a procedure for clock synchronization in the lab frame, as a result of which it can be given any preassigned values within the interval $(0,\theta_{k})$. For instance, in the ultrarelativistic domain, the orientation of the modulation wavefront is always perpendicular to the electron beam velocity (i.e. $\theta_{tilt}=0$) when the evolution of the modulated electron beam is treated using Lorentz coordinates. No physical effects may depend on an arbitrary constant or an arbitrary function.